Please help me solve this equation, 4-3(-6)-(-7)
Thank You.

Found 2 solutions by amarjeeth123, stanbon:
Answer by amarjeeth123(569)

4-3(-6)-(-7)=4+18+7=22+7=29

Answer by stanbon(75887)

4-3(-6)-(-7)
----
4+18+7
----
= 29
==========
Cheers,
Stan H.